# Quantifying Privacy Leakage in Split Inference via Fisher-Approximated Shannon Information Analysis

### v1

@@ Abstract
Split inference (SI) partitions deep neural networks into distributed sub-models, enabling privacy-preserving collaborative learning. Nevertheless, it remains vulnerable to Data Reconstruction Attacks (DRAs), wherein adversaries exploit exposed smashed data to reconstruct raw inputs. Despite extensive research on adversarial attack-defense games, a shortfall remains in the fundamental analysis of privacy risks. This paper establishes a theoretical framework for privacy leakage quantification using information theory, defining it as the adversary’s certainty and deriving both average-case and worst-case error bounds. We introduce Fisher-approximated Shannon information (FSInfo), a novel privacy metric utilizing Fisher Information (FI) for operational privacy leakage computation. We empirically show that our privacy metric correlates well with empirical attacks and investigate some of the factors that affect privacy leakage, namely the data distribution, model size, and overfitting.
Machine Learning, ICML
1 INTRODUCTION
Machine learning as a service (MLaaS) allows users to access machine learning (ML) services by simply uploading the raw data to be inferred. To enable privacy-preserving MLaaS, model partition, a mainstream research line of distributed machine learning, has been widely employed to segment deep neural networks (DNNs) into multiple sub-models and deploy them on different hosts (e.g., mobile phones and edge servers). These hosts conduct collaborative learning by communicating the data derived from the split layer (smashed data), thus hopefully preserving raw data privacy. This learning paradigm has been adopted in numerous model-distributed co-training and co-inference; for example, split/collaborative inference (SI) (Zhang et al., 2021; Li et al., 2021), split learning (SL) (Deng et al., 2023; Yu et al., 2024; Oh et al., 2022), vertical federated learning (VFL) (Chen et al., 2023), pipeline parallelism (Yuan et al., 2023; Ye et al., 2024), and their combinations (Duan et al., 2022). W.l.o.g, we focus on a two-party split inference system (see Figure 1) where the inputs are first fed into the bottom model on the client, and then the intermediate smashed data are transmitted to the top model on the server for further prediction, following Maeng et. al. (Maeng et al., 2023) and Singh et. al. (Singh et al., 2024). It is worth noting that this system represents each client-server pair in a multi-party distributed framework (e.g., VFL) and also corresponds to the forward computation in co-training (e.g., SL).
@@ 1 INTRODUCTION
Despite transmitting only smashed data rather than raw inputs, the SI paradigm remains susceptible to privacy inference attacks. Adversaries can execute Membership Inference Attacks (MIAs) to identify training data participation (Song & Mittal, 2021; Wu et al., 2024), Attribute Inference Attacks (AIAs) to deduce sensitive attributes (e.g., demographic information) (Liu et al., 2022), and Data Reconstruction Attacks (DRAs) to recover original user data (He et al., 2019; Yang et al., 2022; Pasquini et al., 2021). Our research focuses on DRAs due to their heightened severity: successful execution compromises entire datasets rather than isolated attributes, potentially causing substantial damage. As illustrated in Figure 1, the adversary (malicious server) can reconstruct input images after seeing the smashed data, demonstrating the critical nature of this vulnerability.
@@
Privacy leakage quantification (or privacy quantification for short), is an independent concept that transcends the dualistic framework of privacy attack and defense. This critical but less investigated aspect of privacy issues on ML is to develop an appropriate measure (i.e., a privacy metric) of private information leakage from the exposed data (e.g., model parameters, gradients, and smashed data) about the raw input data. In this paper, privacy quantification means assessing the private information exposure from smashed data under DRAs in the SI system of Figure 1. The challenges for this can be presented through the lens of the twofold requirements: i) the privacy metric should be theoretically principled and possess a semantic interpretation aligned with privacy and security settings; and ii) the privacy metric should be quantitative and operationally computable, not just theoretical guidance.
Privacy quantification methods primarily fall into two categories: attack-based and theory-based approaches. Attack-based methods (Wu et al., 2024; Hu et al., 2023; Deng et al., 2023) measure privacy leakage through inference attacks, using success rates as empirical metrics. While operationally feasible, these methods rely heavily on adversary capabilities and provide only theoretical lower bounds without guarantees. Theory-based approaches employ various frameworks, including differential privacy (DP) (Singh et al., 2023), information theory (mutual information; MI) (Arevalo et al., 2024; Noorbakhsh et al., 2024), and hypothesis testing (Fisher information; FI) (Maeng et al., 2023). However, DP faces compatibility issues with nonlinear transformations in split inference, while MI often proves computationally intensive. Although FI offers theoretical rigor and computational efficiency, it focuses solely on mean squared error (MSE), limiting its privacy assessment scope. Consequently, existing methods exhibit inherent limitations in addressing these dual challenges simultaneously.
Contributions. In this paper, we tackle the two challenges by answering the following three questions: i) how to formulate privacy leakage theoretically; ii) how to easily compute the privacy leakage; and iii) what factors may potentially impact privacy leakage. For the first question, we formulate the privacy leakage as the certainty/confidence of the adversary with conditional entropy. Then, for the second question, we employ FI as a tool to directly calculate an approximation of defined privacy leakage, offering a quantitative and operational privacy metric called Fisher-approximated Shannon information (FSInfo) that also considers the adversary’s ability to exploit information. Finally, with FSInfo, we experimentally analyze the determinants of privacy leakage, including data distribution, model width and depth, and overfitting.
2 RELATED WORK
Table 1: Comparisons of the privacy metrics.
Metrics Bound Comp. Adv. Fea. Adv. Err.
Attack
Dcor
MI
FI
FSInfo
@@ 2 RELATED WORK
The privacy metrics (Wagner & Eckhoff, 2018) or privacy quantification, although having been studied decades ago in the areas of smart meter (Kalogridis et al., 2010) and database (Agrawal & Aggarwal, 2001), does not receive much attention in the field of machine learning until recently (Deng et al., 2024).
@@
Current privacy metrics primarily emerge as byproducts of privacy attacks, demonstrating that attacks developed for breaching privacy can also be applied to quantify it (Wu et al., 2024; Hu et al., 2023; Deng et al., 2023; Abuadbba et al., 2020; Pasquini et al., 2021; Liu et al., 2022). For instance, Liu et al. (Liu et al., 2022) developed the ML-Doctor framework, employing four distinct privacy inference attacks to assess private data and model parameter vulnerabilities. However, these empirical approaches, while intuitive, lack theoretical foundations and exhibit adversary-dependent limitations, potentially yielding overly optimistic privacy assessments with weaker adversaries (Deng et al., 2024).
@@
Theoretical approaches offer deeper privacy analysis through advanced mathematical tools. Pan et al. (Pan et al., 2022) quantify privacy leakage via activated neuron counts, modeling it as a linear equation system’s complexity. However, this method primarily applies to fully connected networks (FCNs) with ReLU activations, limiting its effectiveness for nonlinear functions. While differential privacy (DP) remains a prominent DNN privacy framework, traditional local DP proves inadequate for both split inference and reconstruction attacks (Li et al., 2021; Singh et al., 2023; Tan et al., 2024). Singh et al. (Singh et al., 2023) address this limitation by generalizing DP to geodesic on the data manifold, yet DP’s inherent indistinguishability property, while effective against membership inference, demonstrates reduced efficiency and limited efficacy against reconstruction attacks, as empirically shown in (Tan et al., 2024).
Information-theoretic approaches, particularly Shannon mutual MI (Mireshghallah et al., 2020; Noorbakhsh et al., 2024; Arevalo et al., 2024) and distance correlation (Deng et al., 2024; Sun et al., 2022, 2021), quantify privacy leakage by measuring the information shared between $A$ and $B$. Empirical studies demonstrate strong correlations between these metrics and adversarial recovery efficacy, including MSE and semantic similarity (Deng et al., 2024). Researchers have applied these measures to various data pairs: input features and smashed data (Mireshghallah et al., 2020; Noorbakhsh et al., 2024; Deng et al., 2024; Sun et al., 2021), training data and model parameters (Tan et al., 2024), and labels and smashed data (Sun et al., 2022). However, Shannon MI computation remains intractable for high-dimensional data (Arevalo et al., 2024), prompting recent approaches to circumvent direct calculation through noise injection (Tan et al., 2024) or adversarial learning (Noorbakhsh et al., 2024). Moreover, the previous MI-based privacy formulations neglect the characterization of the adversary.
FI metric (Hannun et al., 2021; Guo et al., 2022; Maeng et al., ; Maeng et al., 2023), an easily computable privacy metric with a single backward pass, measures the error of the DRA adversary with theorems in hypothesis testing. However, they consider the mean squared error of the adversary as the privacy definition, which is only a specific part of the privacy.
As shown in Table 1, compared with the existing privacy metrics that can be adapted to the SI settings, FSInfo gives the theoretical lower bound on the adversary’s error (Bound) and is easily computable (Comp.). Also, its privacy leakage formulation considers the adversary’s power or the adversary’s feature of utilizing leaked information (Adv. Fea.) and is proved to have correlation with a general form of the adversary’s error function (Adv. Err).
3 PRELIMINARIES
3.1 Split Inference Pipeline
@@
W.l.o.g, we consider a two-party split inference paradigm between the client $D_{1}$ and the server $D_{2}$ as shown in Figure 1. We focus on the privacy leakage from the smashed data of split inference about the raw input data. $\mathcal{U}=\{(x_{i},y_{i})_{i=1}^{N}\subseteq\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\}$ is a dataset corresponding to intelligent ML services. The two parties learn from $\mathcal{U}$ with a split DNN $f_{\theta}(\cdot)=(f_{\theta 2}\circ f_{\theta 1})(\cdot)=f_{\theta 2}(f_{\theta 1}(\cdot))$, where $D_{1}$ holds the bottom model $f_{\theta 1}$ and $D_{2}$ the top model $f_{\theta 2}$. The split point (SP) $p\in\{1,2,\dots,L\}$ is the last layer of the bottom model $f_{\theta 1}$, where $L$ is the total number of layers of $f_{\theta}$. The procedure starts with that $D_{1}$ takes as inputs a batch of raw data $x^{B}$ and generates smashed data $z^{B}$ as outputs; that is,
$z^{B}=f_{\theta 1}(x^{B}),$ (1)
where $B$ is the batch size. Then, $P_{2}$ forwards $z$ to get the predictions:
$y^{B}=f_{\theta 2}(z^{B}).$ (2)
For simplicity, we use $x$ to denote a batch of raw data if there is no ambiguity.
3.2 Threat Model
@@
We focus on the common setting of DRAs, where the mobile device with private data is benign and the edge server is untrusted.
@@
In a data reconstruction attack (model inversion attack) (Yin et al., 2023; He et al., 2019), the goal of the adversary (edge server) $\mathcal{A}$ is to reconstruct the value of raw data $x$:
$\hat{x}=g_{\phi}(z),$ (3)
where the $g_{\phi}$ is a attack function parameterized by $\phi$, and $\hat{x}$ is the reconstructed input.
@@
As many DRAs rely on different prior knowledge to conduct reconstruction, less assumption about the adversary’s knowledge is preferred for broadening FSInfo’s applicability. Therefore, we do not leave tight requirements on the adversaries’ knowledge by allowing them to know and decide whether to use or not the plain-text smashed data $z$ and all side information. i.e., the auxiliary dataset $\mathcal{V}$ with the same distribution of $\mathcal{U}$, architecture and parameters of bottom model $f_{\theta 1}$. This is compatible with most attacks (He et al., 2019; Yin et al., 2023; Yang et al., 2022) and the same as other privacy metrics (Maeng et al., 2023; Pasquini et al., 2021). For simplicity, we only consider the unbiased attack function $g_{\phi}$, but one can apply our metric to the biased scenarios by using score matching techniques in (Maeng et al., 2023).
4 Privacy Leakage Quantification
4.1 Privacy and Privacy Leakage
The semantic meaning of privacy (information) depends on the diverse attributes of users and application contexts, e.g., culture and law. In contrast, privacy leakage, no matter what the privacy is (e.g., age or facial features), mainly depends on the ML task and threat model, i.e., the specific adversaries and their capabilities. For example, against the DRA adversaries in image classification tasks, privacy leakage is the value of training data $x$. Against MIA adversaries, privacy leakage is the membership of a single data $x$, i.e., the knowledge of whether or not $x$ is used in training data of the model (Singh et al., 2023).
4.2 Privacy Leakage Formulation
In a SI system, we treat the raw input and the corresponding output of bottom model $f_{\theta 1}$ as two multi-dimensional random variables, $X$ and $Z$, characterized by distributions $P(X)$ and $P(Z;X,\theta 1)$. After observing the $Z$, the adversary $\mathcal{A}$ out put the estimation $\hat{X}$ of $X$. This process forms a Markov chain $X\rightarrow Z\rightarrow\hat{X}$.
@@ 4.2 Privacy Leakage Formulation
$\frac{\mathbb{E}_{p(X,\hat{X})}[\|X-\hat{X}\|^{2}]}{d_{x}}\geq\frac{e^{\frac{2}{d_{x}}H(\hat{X}|X)}}{2\pi e}.$ (5)
Theorem 4.2 indicates that the lower bound of the adversary’s error is smaller under a higher privacy leakage.
Strongest Adversary (SA). Definition 4.1 captures the adversary’s uncertainty of $\hat{X}$ given $X$, serving as a naive criterion for measuring the recovery hardness. However, this definition remains agnostic to adversarial power, i.e., how accurate estimation the adversary can achieve. For instance, neural network (NN)-based adversary (He et al., 2019) typically achieve superior reconstruction accuracy compared to maximum likelihood estimation (MLE)-based approaches (He et al., 2019) as its $p(\hat{X})$ is closer to the $p(X)$ (He et al., 2019). To address this limitation, we extend the privacy leakage definition by incorporating adversarial power assumptions. We consider the “strongest” adversary as the one whose estimation $\hat{X}^{s}$ has the least uncertainty among all possible estimators. This approach fundamentally differs from differential privacy’s data-independent worst-case analysis, instead considering the adversarial worst-case across all potential attackers. Here, ”strongest” specifically denotes reconstruction capability rather than knowledge or threat model characteristics.
Following this assumption, we introduce the definition of privacy leakage against the strongest adversary:
Definition 4.3.
@@
We then derive the relation between the privacy leakage definition in Definition 4.3 and the worst-case lower bound of the adversary’s reconstruction error. We adopt the statistical minimax estimation setting (Scarlett & Cevher, 2019), to denote estimation loss as:
$\ell(X,\hat{X})=\Phi(\rho(X,\hat{X})),$ (7)
where $\rho(X,\hat{X})$ is a metric and $\Phi(\cdot)$ is an increasing function maps from $\mathbb{R}^{+}$ to $\mathbb{R}^{+}$. Equation 7 is a general formulation shared among many loss functions. For example, the MSE loss ($\ell(X,\hat{X})=\frac{1}{d_{x}}\|X-\hat{X}\|_{2}^{2}$) used by (Maeng et al., 2023) as the privacy definition clearly takes this form (Scarlett & Cevher, 2019). The minimax risk of the adversary’s estimation is
$\mathcal{M}(\mathcal{X},\ell)$ $=\inf_{\hat{X}\in\mathcal{\hat{X}}}\sup_{X\in\mathcal{X}}\mathbb{E}_{X}[\ell(X,\hat{X})]$ (8)
@@
where $\hat{X}$ are lie in subset $\mathcal{X}$. $\mathcal{M}(\mathcal{X},\ell)$ can be interpreted as the largest estimation error achieved by the strongest DRA adversary in this paper.
Then we can give the lower bound of the adversary’s minimax risk with the variation of Fano inequality (Scarlett & Cevher, 2019) as:
Theorem 4.4.
@@
$\mathcal{M}(\mathcal{X},\ell)\geq\Phi(\frac{\epsilon}{2})(1-\frac{H(\hat{X})-H(\hat{X}|X)+log2}{log(|\mathcal{X}|)}),$ (11)
@@
Due to space limitations, the proof for Theorem 4.2 and Theorem 4.4 is provided in the Appendix. Intuitively, this worst-case lower bound formally characterizes the quantitative relationship between privacy leakage (Definition 4.3) and the adversary’s minimax risk, where $\ell$ represents a general loss function adaptable to specific application scenarios. Specifically, increased privacy leakage $\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}=-H(\hat{X}^{s}|X)$ directly corresponds to a lower minimax risk and less adversarial estimation error.
4.3 Computing Privacy Leakage
Definition 4.3 gives a privacy leakage definition as adversaries’ certainty. However, it can not be used directly for practitioners to evaluate the security level of a split inference framework, as the calculation of conditioned entropy or the MI of arbitrary high-dimensional random variables is a long-standing open problem (Noorbakhsh et al., 2024; Arevalo et al., 2024; Xiao & Devadas, 2023). Therefore, as shown in Figure 2, we translate Fisher information into a Shannon information quantity to estimate the certainty of the “strongest” adversary, i.e. $\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}$. Then we use the Gaussian distribution to get maximum conditional entropy, offering an operational privacy metric, FSInfo.

### v2

@@ Abstract
Split inference (SI) partitions deep neural networks into distributed sub-models, enabling collaborative learning without directly sharing raw data. However, SI remains vulnerable to Data Reconstruction Attacks (DRAs), where adversaries exploit exposed smashed data to recover private inputs. Despite substantial progress in attack–defense methodologies, the fundamental quantification of privacy risks is still underdeveloped. This paper establishes an information-theoretic framework for privacy leakage in SI, defining leakage as the adversary’s certainty and deriving both average-case and worst-case error lower bounds. We further introduce Fisher-approximated Shannon information (FSInfo), a new privacy metric based on Fisher Information (FI) that enables operational and tractable computation of privacy leakage. Building on this metric, we develop FSInfoGuard, a defense mechanism that achieves a strong privacy–utility tradeoff. Our empirical study shows that FSInfo is an effective privacy metric across datasets, models, and defense strengths, providing accurate privacy estimates that support the design of defense methods outperforming existing approaches in both privacy protection and utility preservation. The code is available at https://github.com/SASA-cloud/FSInfo.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I INTRODUCTION
Machine learning as a service (MLaaS) allows users to access machine learning (ML) services by simply uploading the raw data to be inferred. To enable privacy-preserving MLaaS, model partition, a mainstream research line of distributed machine learning, has been widely employed to segment deep neural networks (DNNs) into multiple sub-models and deploy them on different hosts (e.g., mobile phones and edge servers). This learning paradigm has been adopted in numerous model-distributed co-training and co-inference; for example, split/collaborative inference (SI) [zhang2021autodidactic, ICLR21-two-party-label], split learning (SL) [deng2023hsfl, ijcai2024p596, oh2022locfedmix], vertical federated learning (VFL) [chen2023practical], pipeline parallelism [yuan2023pipeedge, ye2024galaxy], and their combinations [duan2022combined]. W.l.o.g, we focus on a two-party split inference system (see Fig. 1) where the inputs are first fed into the bottom model on the client, and then the intermediate smashed data are transmitted to the top model on the server for further prediction, thus hopefully preserving raw data privacy.
@@ I INTRODUCTION
Despite transmitting only smashed data rather than raw inputs, the SI paradigm remains susceptible to privacy inference attacks. Adversaries can execute Membership Inference Attacks (MIAs) to identify training data participation [sec21-evaluation-ml, wu2024quantifying], Attribute Inference Attacks (AIAs) to deduce sensitive attributes (e.g., demographic information) [liu2022ml], and Data Reconstruction Attacks (DRAs) to recover original user data [blackbox, yang2022measuring, pasquini2021unleashing]. Our research focuses on DRAs due to their heightened severity: successful execution compromises entire datasets rather than isolated attributes, potentially causing substantial damage.
@@
Privacy leakage quantification (or privacy quantification for short) is an independent concept that transcends the dualistic framework of privacy attack and defense. This critical but less investigated aspect of privacy issues on ML is to develop an appropriate measure (i.e., a privacy metric) of private information leakage from the exposed data (e.g., model parameters, gradients, and smashed data) about the raw input data. In this paper, privacy quantification means assessing the private information exposure from smashed data under DRAs in the SI system of Fig. 1. Privacy quantification serves a dual purpose: to assess the privacy risks of ML systems for oversight and compliance, and to guide the design of effective defenses by informing architecture and method selection.
Privacy quantification methods in ML primarily fall into two categories: attack-based and theory-based. Attack-based methods [wu2024quantifying, hu2023quantifying, deng2023hsfl] assess privacy leakage via inference attacks, using success rates as empirical metrics. Notably, advances in attack strategies can be directly applied for quantification. While practical, these methods depend heavily on adversary strength and lack theoretical rigor, often underestimating privacy risks. In contrast, theory-based methods use formal theoretical frameworks such as metric differential privacy (metric-DP) [singh2023posthoc], information theory (mutual information; MI) [arevalo2024task, noorbakhsh2024inf2guard], and hypothesis testing (Fisher information; FI) [maeng2023bounding] to infer privacy leakage. However, these approaches are either difficult to compute accurately or overly sensitive to outliers, risking overestimation.
Contributions. To address the above two limitations, in this paper, first, we theoretically define privacy leakage as the adversary’s certainty based on information theory and provide guarantees against DRA adversaries. Second, we employ FI to identify the strongest adversary and derive a lower bound on the defined privacy leakage, yielding a practical, training-free metric: Fisher-approximated Shannon Information (FSInfo). Then, with FSInfo, we deveolop a defense method FSInfoGuard for defending against DRA adversaries without significant performance degradation. Finally, we empirically validate FSInfo as both an effective assessment tool and a design guide for SI systems and defense strategies. Experiments on image, tabular, and text data show that FSInfo correlates well with DRA performance, and defenses guided by FSInfo outperform existing methods in privacy-utility tradeoffs. We also analyze three key design factors—data distribution, model width/depth, and overfitting—to provide practical insights for system development.
In summary, the contributions of our paper are as follows:
  • We formalize privacy leakage against DRAs in SI systems within an information-theoretic framework and derive both average-case and worst-case lower bounds on the adversary’s reconstruction error.
  • We introduce a new privacy metric, FSInfo, which leverages FI to operationally compute the defined leakage. Guided by FSInfo, we design FSInfoGuard, a defense that injects Gaussian noise into the smashed data to mitigate DRAs with minimal performance loss.
  • We conduct extensive experiments to validate the effectiveness of FSInfo and FSInfoGuard and analyze three key factors—dataset characteristics, model size, and overfitting—that influence privacy leakage.
II RELATED WORK
@@
The privacy metrics [CSUR-18-metrics-survey], although having been studied decades ago in the areas of smart meter [kalogridis2010privacy] and database [agrawal2001design], have not received much attention in the field of machine learning until recently [deng2024invmetrics].
@@
Current privacy metrics primarily emerge as byproducts of privacy attacks, demonstrating that attacks developed for breaching privacy can also be applied to quantify it [wu2024quantifying, hu2023quantifying, deng2023hsfl, AsiaCCS20-1dCNN, pasquini2021unleashing, liu2022ml]. For instance, Liu et al. [liu2022ml] developed the ML-Doctor framework, employing four distinct privacy inference attacks to assess private data and model parameter vulnerabilities. However, these empirical approaches, while intuitive, lack theoretical foundations and exhibit adversary-dependent limitations, potentially yielding overly optimistic privacy assessments with weaker adversaries.
@@
Theoretical approaches offer deeper privacy analysis through advanced mathematical tools. Pan et al. [pan2022exploring] quantify privacy leakage via activated neuron counts, modeling it as a linear equation system’s complexity. However, this method primarily applies to fully connected networks (FCNs) with ReLU activations, limiting its effectiveness for nonlinear functions. While differential privacy (DP) remains a prominent DNN privacy framework, traditional local DP proves unsuitable for split inference [ICLR21-two-party-label, singh2023posthoc, sec24-defending-FL-ita]. Singh et al. [singh2023posthoc] address this limitation by generalizing DP definition to geodesic on the data manifold. Yet DP’s inherent indistinguishability property, while effective against membership inference, demonstrates reduced efficiency and limited efficacy against reconstruction attacks, as empirically shown in [sec24-defending-FL-ita].
Information-theoretic approaches, particularly Shannon mutual MI [mireshghallah2020shredder, noorbakhsh2024inf2guard, arevalo2024task] and distance correlation [deng2024invmetrics, sun2022label, sun2021defending], quantify privacy leakage by measuring the information shared between $A$ and $B$. Researchers have applied these measures to various data pairs: input features and smashed data [mireshghallah2020shredder, noorbakhsh2024inf2guard, deng2024invmetrics, sun2021defending], training data and model parameters [sec24-defending-FL-ita], and labels and smashed data [sun2022label]. However, the closed-form information quantities (e.g., Shannon MI) of arbitrarily high-dimensional random variables are still an open problem, prompting recent approaches to circumvent direct MI calculation through noise injection [sec24-defending-FL-ita]. Therefore, in ML, sample-based information quantities estimations have been proposed [cheng2020club]. For example, [mireshghallah2020shredder] uses MI calculation tools based on K-nearest neighbor entropy estimation, and [arevalo2024task, sec24-defending-FL-ita, nips24-club] use CLUB, an upper bound of MI and parameterized variational distributions to quantify privacy leakage. However, these learning-based estimators may fail to approximate reliably when the data dimension increases [cheng2020club]. Also, the tightness of these bounds relies on the choice of variational distribution and the accuracy of trained models, which may lead to approximation inaccuracies.
FI metrics [hannun2021measuring, guo2022bounding, maeng2023bounding], computable with a single backward pass, estimate DRA adversary error based on hypothesis testing theory. However, they are susceptible to outlier activations of smashed data, potentially overestimating privacy leakage.
III PRELIMINARIES
III-A Split Inference Pipeline
@@
W.l.o.g, we consider a two-party split inference paradigm between the client $D_{1}$ and the server $D_{2}$ as shown in Fig. 1. We focus on the privacy leakage from the smashed data of split inference about the raw input data. $\mathcal{U}=\{(x_{i},y_{i})_{i=1}^{N}\subseteq\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\}$ is a dataset corresponding to intelligent ML services. The two parties learn from $\mathcal{U}$ with a split DNN $f_{\theta}(\cdot)=(f_{\theta 2}\circ f_{\theta 1})(\cdot)=f_{\theta 2}(f_{\theta 1}(\cdot))$, where $D_{1}$ holds the bottom model $f_{\theta 1}$ and $D_{2}$ the top model $f_{\theta 2}$. The split point (SP) $p\in\{1,2,\dots,L\}$ is the last layer of the bottom model $f_{\theta 1}$, where $L$ is the total number of layers of $f_{\theta}$. The procedure starts with that $D_{1}$ takes as input the raw data $x$ and generates smashed data $z=f_{\theta 1}(x)$ as output. Then, $P_{2}$ forwards $z$ to get the predictions: $y=f_{\theta 2}(z)$.
III-B Threat Model
@@
We focus on the common setting of DRAs, where the client $D_{1}$ with private data is benign and the server $D_{2}$ is untrusted.
@@
In a data reconstruction attack (model inversion attack) [yin2023ginver, blackbox], the goal of the adversary (edge server) $\mathcal{A}$ is to reconstruct the value of raw data $x$: $\hat{x}=g_{\phi}(z)$, where the $g_{\phi}$ is a attack function parameterized by $\phi$, and $\hat{x}$ is the reconstructed input.
@@
First, as many DRAs rely on different prior knowledge to conduct reconstruction, fewer assumptions about the adversary’s knowledge are preferred for broadening FSInfo’s applicability. Therefore, we do not leave tight requirements on the adversaries’ knowledge by allowing them to know and decide whether to use or not the plain-text smashed data $z$ and all side information. i.e., the auxiliary dataset $\mathcal{V}$ with the same distribution of $\mathcal{U}$, architecture and parameters of bottom model $f_{\theta 1}$. This is compatible with most attacks [blackbox, yin2023ginver, yang2022measuring] and the same as other privacy metrics [maeng2023bounding, pasquini2021unleashing]. Then, for simplicity, we only consider the unbiased attack function $g_{\phi}$, but one can apply our metric to the biased scenarios by using score matching techniques in [maeng2023bounding].
IV Privacy Leakage Quantification
IV-A Privacy and Privacy Leakage
The semantic meaning of privacy (information) depends on the diverse attributes of users and application contexts, e.g., culture and law. In contrast, privacy leakage, no matter what the privacy is (e.g., age or facial features), mainly depends on the ML task and threat model, i.e., the specific adversaries and their capabilities. For example, against the DRA adversaries in image classification tasks, privacy leakage is the value of the input data $x$. Against MIA adversaries, privacy leakage is the membership of a single data $x$, i.e., the knowledge of whether or not $x$ is used in the training data of the model [singh2023posthoc].
IV-B Privacy Leakage Formulation
In a SI system, we treat the raw input and the corresponding output of bottom model $f_{\theta 1}$ as two multi-dimensional random variables, $X$ and $Z$, characterized by distributions $P(X)$ and $P(Z|X,\theta 1)$. After observing the $Z$, the adversary $\mathcal{A}$ out put the estimation $\hat{X}$ of $X$. This process forms a Markov chain $X\rightarrow Z\rightarrow\hat{X}$.
@@ IV-B Privacy Leakage Formulation
$\frac{\mathbb{E}_{p(X,\hat{X})}[\|X-\hat{X}\|^{2}]}{d_{x}}\geq\frac{e^{\frac{2}{d_{x}}H(\hat{X}|X)}}{2\pi e},$ (2)
where the lower bound of the adversary’s error is smaller under a higher privacy leakage. The proof is provided in the Appendix.
Definition IV.1 captures the adversary’s uncertainty of $\hat{X}$ given $X$, serving as a naive criterion for measuring the recovery hardness. However, first, this definition remains agnostic to adversarial power, i.e., how accurate estimation the adversary can achieve. For instance, neural network (NN)-based adversaries typically achieve superior reconstruction accuracy compared to maximum likelihood estimation (MLE)-based approaches as its p(X^)p($\hat${X}) is closer to the $p(X)$ [blackbox]. Second, the corresponding Theorem IV.2 only gives average error over all possible adversaries. But privacy is not an average-case metric [aerni2024evaluations], and it is necessary for the worst-case privacy leakage analysis to avoid damages resulting from the strongest adversary.
Strongest Adversary (SA). To address this limitation, we extend the definition of privacy leakage by incorporating adversaries’ capabilities. We define the “strongest” adversary as the one whose estimate $\hat{X}^{s}$ achieves the highest accuracy among all possible estimators. For example, if the estimate is unbiased, the SA corresponds to the one with the minimum covariance. Note that “strongest” refers to estimation accuracy, not to the adversary’s prior knowledge or its ability to influence the SI process. Importantly, this worst-case analysis is fundamentally different from the data-independent worst-case analysis used in DP.
Following this, we introduce the definition of privacy leakage against the strongest adversary:
Definition IV.3.
@@
We then derive the relation between the privacy leakage definition in Definition IV.3 and the worst-case lower bound of the adversary’s reconstruction error. We adopt the statistical minimax estimation setting [scarlett2019introductory], to denote estimation loss as:
$),\ell(X,\hat{X})=\Phi(\rho(X,\hat{X}$)), (4)
[…]
@@
where $\hat{X}$ are lie in subset $\mathcal{X}$. $\mathcal{M}(\mathcal{X},\ell)$ can be interpreted as the largest estimation error achieved by the SA or the least privacy leakage captured by the SA in this paper.
Then we can give the lower bound of the adversary’s minimax risk with the variation of Fano inequality [scarlett2019introductory] as:
Theorem IV.4.
@@
$\mathcal{M}(\mathcal{X},\ell)\geq\Phi(\frac{\epsilon}{2})(1-\frac{H(\hat{X}^{s})-H(\hat{X}^{s}|X)+log2}{log(|\mathcal{X}|)}),$ (8)
@@
Specifically, $\ell$ represents a general loss function adaptable to specific application scenarios, and increased privacy leakage $\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}=-H(\hat{X}^{s}|X)$ directly corresponds to a lower minimax risk and less adversarial estimation error. The proof for Theorem IV.4 is provided in the Appendix.
IV-C Computing Privacy Leakage
Definition IV.3 gives a privacy leakage definition as adversaries’ certainty. However, it can not be used directly for practitioners to evaluate the security level of a split inference framework, as the calculation of conditioned entropy or the MI of arbitrary high-dimensional random variables is a long-standing open problem [noorbakhsh2024inf2guard, arevalo2024task, xiao2023pac]. Therefore, as shown in Fig. 2, we translate Fisher information into a Shannon information quantity to estimate the certainty of the “strongest” adversary, i.e. $\mathcal{L}^{\mathcal{A}^{s}}_{f_{\theta 1},X}$. Then we use the Gaussian distribution to get maximum conditional entropy, offering an operational privacy metric, FSInfo.